# thanks: Corresponding author

### v3

@@ II.4 A numerical example
Figure 3: The example of classifier with two dimensional vectors. The top left panel is the result of QSN with $c\to\infty$ (with direct comparison of probabilities of each amplitudes). The top-right panel is the result of VQSN with $l=3$ and $c\to\infty$. The bottom-left panel is the result of VQSN with $l=3$ and $c=1000$. The bottom-right is the result of VQSN on tianyan176-2 with $l=3$ and $c=1000$.
In the numerical example, we use $N=32$ two dimensional points which are divided into four classes equally. The training set is generated using scikit-learn [73], and $x$, $y$ coordinates are translated as zenith and azimuth angles of the Bloch sphere, $\theta=(x+1)\pi/2$, $\varphi=(y+1)\pi$. When training, we use ‘Adam’ optimizer [74], and the initialization of the ansatz parameters is randomly sampled from the many-body localized distribution [75], the training is done by using the QuEST toolkit [76]. When $l=3$, the ansatz is trained to a fidelity of $99.4\%$. In terms of quantum control instruction set (QCIS), the circuit for classification is depicted in Fig. 2. Then we scan the Bloch sphere and use QSN and VQSN to determine the classification assignment of each point on the Bloch sphere. The discrimination circuit is also run on tianyan176-2 of Tianyan quantum computing cloud platform, with Zuchongzhi superconducting quantum computing system [77].
@@
III Applications in the search for new physics
@@
III.1 Preparation of the datasets
The datasets are generated using MadGraph5 toolkit[78, 79, 80]. A parton shower is applied using Pythia8 with NNPDF [81, 82], the fast detector simulation is applied using Delphes [83] with either the CMS or muon collider card, depending on the processes under study. The preparation of the datasets are applied using MLAnalysis [84].
After the collision events are generated, a set of observables are picked to form the feature space, an event can be mapped into this feature space as a vector $\vec{v}$. Assuming $D$ observables are picked, then $\vec{v}$ is a $D$-dimensional vector. The z-score standardization [85] is applied, $x^{j}_{m}=\left(v^{j}_{m}-\bar{v^{j}}\right)/\epsilon^{j}$, where $v^{j}_{m}$ is the $j$-th component ($10\leq j\leq D-$ 1) of $\vec{v}_{m}$ for the $m$-th collision event, $\bar{v^{j}}$ and $\epsilon^{j}$ are mean values and standard deviations of the $j$-th component of $\vec{v}$ over the training dataset.
To map the above vector to a quantum state, we use amplitude encoding,
@@ III.1 Preparation of the datasets
where $d$ is the least power of $2$ larger than $d^{\prime}/2$, for $x^{j}_{m}$ with $j>D-1$, $x^{j=d^{\prime}}_{m}=1$ and $x^{j>d^{\prime}}_{m}=0$ are used. The encoding in Eq. (2) is for convenience. This study has not addressed further improvements such as encoding the vectors in the dataset into a quantum state using a nonlinear approach [86]. It is also possible to utilize the vast Hilbert space of quantum computers to separate the signal from the background more when mapping to the quantum feature space [87].
III.2 Event selection strategy
The search for NP is binary classification, therefore we change the last step of (V)QSN in Sec. II.3. Denoting $\psi_{n,{\rm SM}}$ and $\psi_{n,{\rm NP}}$ as states for the SM and NP events, respectively, the state to be trained is arranged as,
@@ III.2 Event selection strategy
The number of $|0\rangle_{b_{1}}$ outcome is denoted as $a$ and $\hat{a}=a/(rc)$ is used as a score to indicate the probability such how likely $\phi$ is from the NP, where $r\times c$ is the number of times $\langle\phi|\Psi\rangle$ is successfully built. And after $\hat{a}$ is obtained, we select events with $\hat{a}>\hat{a}_{th}$ where $\hat{a}_{th}$ is a tunable threshold that maximizes the signal significance which is defined as [88, 89],
@@
For comparison with a classical algorithm, we also introduce the classical KNN algorithm which is already used in the search for NP [22]. We use a modified KNN (denoted as ‘t-KNN’ where ‘t’ stands for tunable) with the following strategy for event selection.
@@
For the (t-)KNN algorithm, we only incorporate the cost needed to compute the distances, so the computational complexity (cc) can be estimated as $\mathcal{O}(d^{\prime}NM)$ for $M$ test events. In the context of quantum computers, the number of quantum gates is typically employed as a means of assessing the cc. Even if the dataset is classical, the cc for VQSN is $n_{t}+\mathcal{O}(c^{\prime}M(n_{ans}+d)\sqrt{d})$, where $n_{t}$ is the cost to train the ansatz which is a constant in the view of test dataset, $n_{ans}$ is the cc of the ansatz, the cc of $U_{|\phi\rangle}^{\dagger}$ is $\mathcal{O}(d)$, and $c=c^{\prime}/\sqrt{d}$ where $1/\sqrt{d}$ comes from the success rate to create the state $\langle\phi|\Psi(\alpha_{i})\rangle$ when amplitude amplification is used and the number of times of measurement on $a$ is required to be a fixed number $c^{\prime}$. Note that, $n_{t}$ depends on $N$ and $d$, but not on $M$. Unlike $M$ which grows with the luminosity, $N$ and $d$ remain effectively constant in practical applications, as a result, for practical applications it can be expected that $M\gg N$. By omitting $n_{t}$ because $M\gg N$, and by the fact that $d$ is at the same order of $D$, it can be estimated that the VQSN outperforms (t-)KNN even with a classical dataset when $c(n_{ans}+d)<N\sqrt{d}$. From the principle of the algorithm, the larger $N$ and $d$, the better the classification, as long as $M\gg N$ is kept, this means that ideally VQSN can outperform (t-)KNN even dealing with a classical dataset.
III.3 Search for the gluon quartic gauge couplings at a muon collider
III.3.1 The contribution of the gluon quartic gauge coupling at a muon collider
@@
The gluon quartic gauge couplings (gQGCs) emerge from the Born-Infeld (BI) extension of the SM, which was originally driven by the idea that there should be an upper limit on the strength of the electromagnetic field [90]. It has been demonstrated that the BI model is also associated with M-theory-inspired models [91, 92, 93, 94, 95]. In the SMEFT, the operators contributing to gQGCs appear at dimension-8. As the importance of the dimension-8 operators in the SMEFT has been realized [96, 97, 98, 99, 100, 101], more and more phenomenological studies have been devoted to the dimension-8 operators in recent years [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123]. The case of gQGC is among the operators of dimension-8 in the SMEFT that have received attention. We focus on the $O_{gT,0}$ operator,
@@
where ${G_{\mu\nu}^{a}}$ is gluon field strengths, ${W_{\mu\nu}^{i}}$ denote electroweak field strengths, and $M_{0}$ is the NP scale. The expected constraints is $M_{0}\geq 1040\;{\rm MeV}$ at the LHC with the center of mass (c.m.) energy $\sqrt{s}=13\;{\rm TeV}$, and luminosity $L=36.7\;{\rm fb}^{-1}$ obtained by using the process $gg\to\gamma\gamma$ [91]. The combined sensitivities of the $Z\gamma$ and $\gamma\gamma$ channels at the LHC with $\sqrt{s}=13\;{\rm TeV}$ and $L=137\;{\rm fb}^{-1}$ are about three times of the above [92]. For convenience, we define $f_{0}\equiv 1/(16M_{0}^{4})$.
At muon colliders, the gQGCs contributes to the process $\mu^{+}\mu^{-}\to\nu\bar{\nu}jj$ [124]. The Feynman diagrams are shown in Fig. 4. This process is chosen as one of the processes to validate the VQSN since it is convenient to handle due to the absence of interference between the SM and NP.
III.3.2 Training of the ansatz
@@
$2^{17}+2^{17}$ events are selected from the SM and NP events to form the training set. Since the $XY2M(\theta)=R_{z}(\theta)R_{x}(-\pi/2)R_{z}(-\theta)$ and $XY2P(\theta)=R_{z}(\theta)R_{x}(\pi/2)R_{z}(-\theta)$ are supported gates on tianyan176-2, the single qubit layer of the ansatz are chosen to be $XY2M(\theta_{1})XY2P(\theta_{2})$ on each qubit. The entanglement layer is chosen to be pairwise. The ansatz is chosen to have $l=3$ entanglement layers, with $N=2^{18}$ and $d=4$, $n_{q}=20$ qubits are used. The ansatz is trained using QuEST, the fidelity reaches $78.5\%$.
@@ III.3.3 Numerical results
where $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ are the cross-section of the SM and NP after requiring at least two jets, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ is obtained when $f_{0}=\tilde{f}_{0}$. At $\sqrt{s}=3\;{\rm TeV}$ and $\tilde{f}_{0}=1\;{\rm TeV}^{-4}$, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}=722.9\;{\rm fb}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}=3.21\;{\rm fb}$ [124]. By choosing proper $\hat{a}_{th}$ and $\hat{b}_{th}$, the expected constraints on $|f_{0}|$ at $\sqrt{s}=3\;{\rm TeV}$ and $L=1\;{\rm ab}^{-1}$ (where $L$ is the luminosity chosen according to the design of the muon colliders [125, 126]) are listed in Table 2. The expected constraints on $|f_{0}|$ by using a traditional event selection strategy at $\mathcal{S}_{start}=2$, $3$ and $5$ are $<0.155$, $<0.191$ and $0.248\;{\rm TeV}^{-4}$, respectively. The dataset in this paper is in fact a subset (about $28\%$) of the one used in Ref. [124], it can be seen that the VQSN can reach a same order of magnitude, while t-KNN outperforms the traditional method.
@@
It has been shown that, the randomness in measurement contributes to the primary of the errors. In this subsection, $c=1000$ is used. When $c$ is a constant, the accuracy of $\hat{a}$ is affected by $r$, the success rates of building the states $\langle\phi|\Psi_{\rm train}\rangle$ and $\langle\phi|\Psi(\alpha_{i})\rangle$ (where $\phi$ is the state for a test vector). Fig. 8 illustrates the normalized distributions of success rates where $\phi$ are the signal and background events in the validation dataset. The background and signal events are considered separately because the majority of the dataset from experiments will be the latter. It can be observed that the success rate is typically higher for background events than for signal events. This is due to the more centralized distribution of background events since the success rate is $\sum_{n}^{N}\left|\langle\phi|\psi_{n}\rangle\right|^{2}/N$ for QSN. Furthermore, it was found that the ansatz not only did not reduce but rather enhanced the success rates for the background events. This further reduces the cc of VQSN because for a fixed $c^{\prime}$, a smaller $c$ is needed.
The cc of VQSN and its classical counterpart, i.e. (t-)KNN is compared. For one test event, $168$ gates need to be applied in terms of QCIS, therefore, when $c=1000$, the cc is estimated as $168000$. For (t-)KNN, the cc is for one test event is $1.57\times 10^{6}$ because the training set contains $2^{18}$ $6$-dimensional vectors. It can be seen that, even with a classical dataset, the cc of VQSN can be one order of magnitude smaller.
III.4 Search for the gluon quartic gauge couplings at the LHC
III.4.1 The gluon quartic gauge couplings and the process to be considered

### v4

@@ 2.5 The numerical result
Figure 6: The example of classifier with two dimensional points. The top left panel is the result of QSN with $c\to\infty$ (with direct comparison of probabilities of each amplitudes). The top-right panel is the result of VQSN with $l=3$ and $c\to\infty$. The bottom-left panel is the result of VQSN with $l=3$ and $c=1000$. The bottom-right is the result of VQSN on tianyan176-2 with $l=3$ and $c=1000$.
In the numerical example, we use $N=32$ two dimensional points which are divided into four classes equally. The training set is generated using scikit-learn Pedregosa:2011ork , and $x$, $y$ coordinates are translated as zenith and azimuth angles of the Bloch sphere, $\theta=(x+1)\pi/2$, $\varphi=(y+1)\pi$. When training, we use ‘Adam’ optimizer Kingma:2014vow , and the initialization of the ansatz parameters is randomly sampled from the many-body localized distribution Park:2024rim , the training is done by using the QuEST toolkit Jones:2019knd . When $l=3$, the ansatz is trained to a fidelity of $99.4\%$. In terms of quantum control instruction set (QCIS), the circuit for classification is depicted in Fig. 5. Then we scan the Bloch sphere and use QSN and VQSN to determine the classification assignment of each point on the Bloch sphere. The discrimination circuit is also run on tianyan176-2 of Tianyan quantum computing cloud platform, with Zuchongzhi superconducting quantum computing system Zhu:2021gkn .
@@
On the tianyan176-2 quantum processor, the average two-qubit gate error is $1.6\%$, the average single-qubit gate error is $0.28\%$, and the average readout error is $3.4\%$. The median decoherence times are $T_{1}=22.89\mu s$ and $T_{2}=2.94\mu s$. If we consider only quantum state fidelity and neglecting single-qubit gate errors, the probability of obtaining a correct result can be estimated as $(1-0.016)^{L}\times(1-0.034)^{n_{m}}$, where $L=3$ is the circuit depth in the sense of two-qubit gates, and $n_{m}$ is the number of qubits measured. This yields an estimated accuracy of $85.9\%$.
However, as observed in Fig. 6, in the absence of quantum noise and with a finite number of measurement shots $c$, mis-tagged points occur near decision boundaries. When real quantum noise is introduced, no significant discrepancies attributable to quantum errors were detected. To evaluate the impact of noise, we treat the result with $l=3$ and $c\to\infty$ as the ground truth. The mis-tag rate is approximately $11\%$ for the quantum simulator (due solely to finite measurements) and about $15\%$ on the real quantum device. This indicates that the majority of mis-tags are due to finite sampling rather than quantum noises.
The noise tolerance of the measurement outcomes depends on both the distribution of the training dataset and the specific test point in question. For instance, if the training set contains only two points, a signal event $\psi_{sig}$ and a background event $\psi_{bg}$, then a more precise measurement is required when $\langle\phi|\psi_{sig}\rangle$ is close to $\langle\phi|\psi_{bg}\rangle$ (i.e., when the test point $\phi$ lies near the decision boundary). In contrast, when the two inner products differ significantly, the measurement outcome is robust to the noise.
Kullback-Leibler (KL) divergence can be used to uncover errors that are masked by this robustness. Denoting $P_{i}$ as the probability of outcome class $i$, KL divergence is defined as $KL=H(P^{truth},P^{test})-H(P^{truth},P^{truth})$, where $H$ denotes cross entropy, $H(P^{1},P^{2})=-\sum_{i}P_{i}^{1}\log(P_{i}^{2})$. The average KL divergence over all test points is $0.003$ for the simulator and $0.026$ for the real quantum computer. Unlike the mis-tag rate, the KL divergence significantly increases for the real quantum computer. That means, although noise alters the probabilities of outcomes, these changes are generally insufficient to alter the most probable class assignment, which is the reason of the robustness of the method.
3 Applications in the search for new physics
@@
3.1 Preparation of the datasets
The datasets are generated using MadGraph5 toolkitAlwall:2014hca ; Christensen:2008py ; Degrande:2011ua . A parton shower is applied using Pythia8 with NNPDF Sjostrand:2014zea ; Ball:2013hta , the fast detector simulation is applied using Delphes deFavereau:2013fsa with either the CMS or muon collider card, depending on the processes under study. The preparation of the datasets are applied using MLAnalysis Guo:2023nfu .
After the collision events are generated, a set of observables are picked to form the feature space, an event can be mapped into this feature space as a vector $\vec{v}$. Assuming $D$ observables are picked, then $\vec{v}$ is a $D$-dimensional vector. The z-score standardization Donoho_2004 is applied, $x^{j}_{m}=\left(v^{j}_{m}-\bar{v^{j}}\right)/\epsilon^{j}$, where $v^{j}_{m}$ is the $j$-th component ($0\leq j\leq D-1$) of $\vec{v}_{m}$ for the $m$-th collision event, $\bar{v^{j}}$ and $\epsilon^{j}$ are mean values and standard deviations of the $j$-th component of $\vec{v}$ over the training dataset. $\vec{x}$ is $\vec{v}$ after z-score standardization.
In the search for NP, the data to be dealt with are the collision events. The idea of this work is to decide whether some unknown events are neighbors of the SM events, or the NP events. In the case of HEP, there are different methods to define the distance between two events, which can have an impact on the results of event selection strategies Komiske:2019fks . In (V)QSN, we use the overlap of two quantum states to calculate the distance, therefore the definition of the distance depends on the method to encode vector $\vec{x}$ into a Hilbert space. There exist various methods for mapping vectors into a Hilbert space. Among these, the most straightforward approach is amplitude encoding. While different schemes may impact the efficiency of event selection, we employ the most straightforward amplitude encoding method for brevity and to maintain focus on (V)QSN, leaving the optimization of the encoding scheme for future work. For example, this study has not addressed improvements such as encoding the vectors in the dataset into a quantum state using a nonlinear approach Havlicek:2018nqz . It is also possible to utilize the vast Hilbert space of quantum computers to separate the signal from the background more when mapping to the quantum feature space Lloyd:2020eeh . The vector $\vec{x}$ is map to the quantum state,
@@ 3.1 Preparation of the datasets
where $d$ is the least power of $2$ larger than $d^{\prime}/2$, for $x^{j}_{m}$ with j>D1j>D-1, $x^{j=d^{\prime}}_{m}=1$ and $x^{j>d^{\prime}}_{m}=0$ are used. The amplitude encoding is also implemented using the uniform rotation gates.
3.2 Event selection strategy
Both (V)QSN and (t-)KNN fundamentally remain traditional cut-and-count approaches. Whereas conventional event selection strategies apply cuts directly to observables in collision events, (V)QSN and (t-)KNN instead target anomaly scores $\hat{a}$ and $\hat{b}$. Essentially, $\hat{a}$ and $\hat{b}$ are themselves observables generated from final-state observable information. The efficacy of these methods hinges critically on using events with SM kinematic features and events exhibiting NP kinematic features (i.e., the events from NP squared matrix element) during training.
During the event selection stage, interference terms must be taken into account, the search for NP is no longer a classification problem. Similar as traditional cut-and-count schemes, the differential cross-section can now be regarded as a function of the anomaly scores $\hat{a}$ or $\hat{b}$. Selecting events is equivalent to adjusting the integration limits of the phase space integral. Therefore, at their core, (V)QSN and (t-)KNN utilize kinematic information from final-state observables to suppress the cross-section of events ‘appear SM-like’ while retaining cross-section that ’appear non-SM-like’, thereby enhancing signal significance. In other words, we use the difference between the total cross-sections with and without NP after cut to ‘discover’ NP or set constraints on the parameters of NP.
Because only two classes are involved, we change the last step of (V)QSN in Sec. 2.3. Denoting $\psi_{n,{\rm SM}}$ and $\psi_{n,{\rm NP}}$ as states for the SM and NP events, respectively, the state to be trained is arranged as,
@@ 3.2 Event selection strategy
The number of $|0\rangle_{b_{1}}$ outcome is denoted as $a$ and $\hat{a}=a/(rc)$ is used as a score to indicate the probability such how likely $\phi$ is from the NP, where $r\times c$ is the number of times $\langle\phi|\Psi\rangle$ is successfully built. And after $\hat{a}$ is obtained, we select events with $\hat{a}>\hat{a}_{th}$ where $\hat{a}_{th}$ is a tunable threshold that maximizes the signal significance which is defined as ParticleDataGroup:2020ssz ; Cowan:2010js ,
@@
For comparison with a classical algorithm, we also introduce the classical KNN algorithm which is already used in the search for NP DeSimone:2018efk . We use a modified KNN (denoted as ‘t-KNN’ where ‘t’ stands for tunable) with the following strategy for event selection.
@@
For the (t-)KNN algorithm, we only incorporate the cost needed to compute the distances, so the computational complexity (cc) can be estimated as $\mathcal{O}(d^{\prime}NM)$ for $M$ test events. In the context of quantum computers, the number of quantum gates is typically employed as a means of assessing the cc. Even if the dataset is classical, the cc for VQSN is $n_{t}+\mathcal{O}(c^{\prime}M(n_{ans}+d)\sqrt{d})$, where $n_{t}$ is the cost to train the ansatz which is a constant in the view of test dataset, $n_{ans}$ is the cc of the ansatz, the cc of $U_{|\phi\rangle}^{\dagger}$ is $\mathcal{O}(d)$, and $c=c^{\prime}/\sqrt{d}$ where $1/\sqrt{d}$ comes from the success rate to create the state $\langle\phi|\Psi(\alpha_{i})\rangle$ when amplitude amplification is used and the number of times of measurement on $a$ is required to be a fixed number $c^{\prime}$. Note that, $n_{t}$ depends on $N$ and $d$, but not on $M$. Unlike $M$ which grows with the luminosity, $N$ and $d$ remain effectively constant in practical applications, as a result, for practical applications it can be expected that $M\gg N$. In other words, $n_{t}$ is independent of the luminosity of future colliders. The ansatz can be viewed as a compressed representation of the training set. Its complexity should be lower than that of the full dataset; otherwise, using an ansatz would offer no advantage. Thus, for each selection operation, the ansatz provides computational savings compared to ansatz-free methods, until the initial training cost is amortized. As $M$ grows (i.e., as the luminosity of future colliders increases), the advantage of VQSN over (t-)kNN becomes more pronounced. By omitting $n_{t}$ because $M\gg N$, and by the fact that $d$ is at the same order of $D$, it can be estimated that the VQSN outperforms (t-)KNN even with a classical dataset when $c(n_{ans}+d)<N\sqrt{d}$. From the principle of the algorithm, the larger $N$ and $d$, the better the classification, as long as $M\gg N$ is kept, this means that ideally VQSN can outperform (t-)KNN in the sense of cc even dealing with a classical dataset.
3.3 Search for the gluon quartic gauge couplings at a muon collider
3.3.1 The contribution of the gluon quartic gauge coupling at a muon collider
@@
The gluon quartic gauge couplings (gQGCs) emerge from the Born-Infeld (BI) extension of the SM, which was originally driven by the idea that there should be an upper limit on the strength of the electromagnetic field Born:1934gh . It has been demonstrated that the BI model is also associated with M-theory-inspired models Ellis:2018cos ; Ellis:2021dfa ; Fradkin:1985qd ; Tseytlin:1999dj ; Cheung:2018oki . In the SMEFT, the operators contributing to gQGCs appear at dimension-8. As the importance of the dimension-8 operators in the SMEFT has been realized Green:2016trm ; Zhang:2020jyn ; Murphy:2020rsh ; Li:2020gnx ; Anders:2018oin ; Henning:2015alf , more and more phenomenological studies have been devoted to the dimension-8 operators in recent years Zhang:2018shp ; Bi:2019phv ; ATLAS:2017vqm ; CMS:2017rin ; CMS:2020ioi ; CMS:2016gct ; CMS:2017zmo ; CMS:2018ccg ; ATLAS:2018mxa ; CMS:2019uys ; CMS:2016rtz ; CMS:2017fhs ; Ellis:2023ucy ; Spor:2022hhn ; Spor:2022zob ; Yilmaz:2021ule ; Ellis:2020ljj ; Senol:2019swu ; Yilmaz:2019cue ; Ellis:2019zex ; Jahedi:2022duc ; Jahedi:2023myu . The case of gQGC is among the operators of dimension-8 in the SMEFT that have received attention. We focus on the $O_{gT,0}$ operator,
@@
where $G_{\mu\nu}^{a}$} is gluon field strengths, $W_{\mu\nu}^{i}$} denote electroweak field strengths, and $M_{0}$ is the NP scale. The expected constraints is $M_{0}\geq 1040\;{\rm MeV}$ at the LHC with the center of mass (c.m.) energy $\sqrt{s}=13\;{\rm TeV}$, and luminosity $L=36.7\;{\rm fb}^{-1}$ obtained by using the process $gg\to\gamma\gamma$ Ellis:2018cos . The combined sensitivities of the $Z\gamma$ and $\gamma\gamma$ channels at the LHC with $\sqrt{s}=13\;{\rm TeV}$ and $L=137\;{\rm fb}^{-1}$ are about three times of the above Ellis:2021dfa . For convenience, we define $f_{0}\equiv 1/(16M_{0}^{4})$.
At muon colliders, the gQGCs contributes to the process $\mu^{+}\mu^{-}\to\nu\bar{\nu}jj$ Yang:2023gos . The Feynman diagrams are shown in Fig. 7. This process is chosen as one of the processes to validate the VQSN since it is convenient to handle due to the absence of interference between the SM and NP.
3.3.2 Training of the ansatz
@@
$2^{17}+2^{17}$ events are selected from the SM and NP events to form the training set. Since the $XY2M(\theta)=R_{z}(\theta)R_{x}(-\pi/2)R_{z}(-\theta)$ and $XY2P(\theta)=R_{z}(\theta)R_{x}(\pi/2)R_{z}(-\theta)$ are supported gates on tianyan176-2, the single-qubit layer of the ansatz are chosen to be $XY2M(\theta_{1})XY2P(\theta_{2})$ on each qubit. The entanglement layer is chosen to be pairwise. The ansatz is chosen to have $l=3$ entanglement layers, with $N=2^{18}$ and $d=4$, $n_{q}=20$ qubits are used. The ansatz is trained using QuEST, the fidelity reaches $78.5\%$.
@@ 3.3.3 Numerical results
where $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ are the cross-section of the SM and NP after requiring at least two jets, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ is obtained when $f_{0}=\tilde{f}_{0}$. At $\sqrt{s}=3\;{\rm TeV}$ and $\tilde{f}_{0}=1\;{\rm TeV}^{-4}$, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}=722.9\;{\rm fb}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}=3.21\;{\rm fb}$ Yang:2023gos . By choosing proper $\hat{a}_{th}$ and $\hat{b}_{th}$, the expected constraints on $|f_{0}|$ at $\sqrt{s}=3\;{\rm TeV}$ and $L=1\;{\rm ab}^{-1}$ (where $L$ is the luminosity chosen according to the design of the muon colliders Black:2022cth ; Accettura:2023ked ) are listed in Table 2. The expected constraints on $|f_{0}|$ by using a traditional event selection strategy at $\mathcal{S}_{start}=2$, $3$ and $5$ are <0.155<0.155, <0.191<0.191 and $0.248\;{\rm TeV}^{-4}$, respectively. The dataset in this paper is in fact a subset (about $28\%$) of the one used in Ref. Yang:2023gos , it can be seen that the VQSN can reach a same order of magnitude, while t-KNN outperforms the traditional method.
@@
It has been shown that, the randomness in measurement contributes to the primary of the errors. In this subsection, $c=1000$ is used. When cc is a constant, the accuracy of $\hat{a}$ is affected by $r$, the success rates of building the states $\langle\phi|\Psi_{\rm train}\rangle$ and $\langle\phi|\Psi(\alpha_{i})\rangle$ (where $\phi$ is the state for a test vector). Fig. 11 illustrates the normalized distributions of success rates where $\phi$ are the signal and background events in the validation dataset. The background and signal events are considered separately because the majority of the dataset from experiments will be the latter. It can be observed that the success rate is typically higher for background events than for signal events. This is due to the more centralized distribution of background events since the success rate is $\sum_{n}^{N}\left|\langle\phi|\psi_{n}\rangle\right|^{2}/N$ for QSN. Furthermore, it was found that the ansatz not only did not reduce but rather enhanced the success rates for the background events. This further reduces the cc of VQSN because for a fixed $c^{\prime}$, a smaller cc is needed. However, this improvement depends on the training data distribution, the ansatz structure, and the data encoding method, and should not be considered a universal result. The ansatz serves as an approximation and simplification of the training set. When the ansatz perfectly matches the training dataset, the state preparation success rates should be identical.
The cc of VQSN and its classical counterpart, i.e. (t-)KNN is compared. For one test event, $168$ gates need to be applied in terms of QCIS, therefore, when $c=1000$, the cc is estimated as $168000$. For (t-)KNN, the cc is for one test event is $1.57\times 10^{6}$ because the training set contains $2^{18}$ 66-dimensional vectors. It can be seen that, even with a classical dataset, the cc of VQSN can be one order of magnitude smaller.
3.4 Search for the gluon quartic gauge couplings at the LHC
3.4.1 The gluon quartic gauge couplings and the process to be considered